OpenStudy (anonymous):

Did I do this Algebra problem correctly? 7^13/7^3

OpenStudy (anonymous):

Simplify. Assume that no denominator is equal to 0. I got 1^10, correct?

hero (hero):

Not exactly

hero (hero):

Apply this property of Exponents: \[\frac{n^a}{n^b} = n^{a - b}\]

OpenStudy (anonymous):

Ah, 7^10?

hero (hero):

Good job

OpenStudy (anonymous):

Oh thank you! Egads, I thought this problem was MUCH harder than it was in reality. Ha ha! Thank you! Thank you so much! :P

OpenStudy (anonymous):

I'm a little embarrassed, to be honest. :)

OpenStudy (anonymous):

Do you think you could explain this? (3a/a^2)^-2

hero (hero):

To see it a little more clearly: 

hero (hero):

You cancel factors of 1 leaving just \(7^{10}\)

OpenStudy (anonymous):

Alright, thanks, I got it. Do you think you could explain the next problem I gave you, please? :)

hero (hero):

\[\left (\large \frac{3a}{a^2} \right)^{-2}\]

OpenStudy (anonymous):

Yes

hero (hero):

Basically, the negative exponent you see has nothing to do with negative

OpenStudy (anonymous):

That's misleading, but okay.

hero (hero):

In reality, it means "inverse"

hero (hero):

It means the `INVERSE` of \[\left (\large \frac{3a}{a^2} \right)^{2}\]

hero (hero):

`The inverse of a number is simply one over that number:` \[a^{-2} = \frac{1}{a^2}\]

hero (hero):

The inverse of "a to the negative 2nd power" equals "1 divided by a to the second power".

hero (hero):

Now, here's what you do with your problem: Rewrite it like this: \[1 \div \left (\large \frac{3a}{a^2} \right)^{2}\]

hero (hero):

And then simplify the \[\left (\large \frac{3a}{a^2} \right)^{2}\] part

OpenStudy (anonymous):

3a^2/a^4

hero (hero):

You should end up with this: \[\large 1 \div \large \frac{9a^2}{a^4} \]

OpenStudy (anonymous):

Ah

hero (hero):

Use the reciprocal rule to rewrite it like this: \[ \large 1 \times \large \frac{a^4}{9a^2} \]

OpenStudy (anonymous):

So.. the same thing. :P

hero (hero):

No, NOT the same

OpenStudy (anonymous):

1/9a^2!

hero (hero):

What did you do with the \(a^4\)?

hero (hero):

Let me finish explaining

OpenStudy (anonymous):

Ok

hero (hero):

Next you use the exponents rule I showed you earlier to re-write it like this: \[\large \frac{a^{4 -2}}{9}\]

OpenStudy (anonymous):

Ah! a^2/9

hero (hero):

I hope I didn't lose you

OpenStudy (anonymous):

Not at all! Although, would you mind staying a bit to see if I can understand another problem that's similar? :)

OpenStudy (anonymous):

Actually, it's more similar to the first problem I gave you, however, this is different.. 12/c^-4d^8

OpenStudy (anonymous):

I think it's 12/c^4d^8, correct?

hero (hero):

If I were you I would rewrite it like this to avoid confusion: \[12 \div (c^{-4} \times d^8)\]

hero (hero):

This way, you can deal with the \(c^{-4}\) separately

hero (hero):

Or you could just leave it as \[\frac{12}{c^{-4}d^8}\] and deal with it that way

hero (hero):

Now here's the thing. If the negative exponent is in the denominator, move it to the numerator and remove the exponent. That way, it will become: \[\frac{12c^4}{d^8}\]

hero (hero):

Are you lost yet?

OpenStudy (anonymous):

A bit. Ha ha! I presume that's the final answer?

hero (hero):

Yes. If you follow exactly as I have explained to you, you won't have any problems.

OpenStudy (anonymous):

Alright! Now, another similar but different problem :P Simplify the expression 6k^-6b^6

hero (hero):

Okay, you should be able to follow what I've shown you to figure this one out